10. Tom's uncle owns a triangular piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one-half of the longest side. The second longest side is 2 yards shorter than the longest side. What is the length of each side?
Answer by checkley77(12844)

S=30+L/2
M=L-2
S+M+L=378
(30+L/2)+(L-2)+L=378
30+L/2+L-2+L=378
L/2+2L=378-30+2
2.5L=350
L=350/2.5
L=140 YDS. FOR THE LONG SIDE
S=30+140/2
S=3+70
S=100 YDS. FOR THE SHORT SIDE.
M=140-2
M=138 YDS. FOR THE MIDDLE SIDE.
PROOF:
140+138+100=378
378=378